You have been hired as a financial advisor to Steph Curry. He has received two offers for playing professional basketball and wants to select the best offer, based on considerations of money only. Offer A is an offer for $25 m a year for four years. Offer B is an offer of $20 m a year for five years. What is your advice, if the interest rate is 10% per year?
Answer #1

Total cash flow in offer a is 25*4=100 million and in offer b is 20*5=100 million

So, total cash flow is same in both the offers. Hence, without any calculation we see that Offer A is better than Offer B because cash flows are being realised earlier (25 versus 20)

Using calculations:

Present Value of Offer A=25/10%*(1-1/1.1^4)=79.24663616 million

Present Value of Offer B=20/10%*(1-1/1.1^5)=75.81573539 million

So, Offer A is better than Offer B